# Address-addressable read AI artificial intelligence

### v1

@@ Abstract
The graph database (GDB) is an increasingly common storage model for data involving relationships between entries. Beyond its widespread usage in database industries, the advantages of GDBs indicate a strong potential in constructing symbolic artificial intelligences and retrieval-augmented generation (RAG), where knowledge of data inter-relationships takes a critical role in implementation. However, current GDB models are not optimised for hardware acceleration, leading to bottlenecks in storage capacity and computational efficiency. In this paper, we propose a hardware-friendly GDB model, called Views. We show its data structure and organisation tailored for efficient storage and retrieval of graph data and demonstrate its equivalence to represent traditional graph representations. We further demonstrate its symbolic processing abilities in semantic reasoning and cognitive modelling with practical examples and provide a short perspective on future developments.
@@
Keywords – graph database, data structure, near-memory computing, semantic processing
@@ 1 Introduction
Because of these factors, building “knowledge graphs” is a key topic in GDBs [1]. These are graphs used to store and convey real-world knowledge. Because of their explicit semantics, knowledge graphs have been widely used as databases in symbolic cognitive computing [2]. Knowledge graphs facilitate access to stored symbols for cognitive systems and offer insights into their processes [3, 4]. They are also increasingly important as knowledge bases for large language models in RAG pipelines to improve large language model accuracy and reduce hallucinations [5].
@@
Functionally, the critical operation in GDB is querying. For example, the query of “find me all films which Tom Hanks has acted in” followed by “who is Tom Hanks”, as anyone might do in any film database. Multiple querying languages and formal frameworks have been proposed over the years for query efficiency improvements, such as GQL, SPARQL, Cypher, etc. However, such high-level querying languages are built upon but not necessarily optimised for existing computing facilities. In the domain of symbolic cognitive computing, where operations require a high volume of content search, the discreteness of concepts presents a bottleneck in efficient processing and scaling, which GDBs in general also struggle with due to internal data multi-modality and inter-database operability issues [6, 7, 8]. To address these problems, contemporary solutions such as GPU clusters have been employed to process GDB workloads, whereas high costs in power and manufacture hinder their mass usage [9].
@@
In this direction, Pratt developed the initial hierarchical graph model of recursive labelling for semantics of programming languages [10], and Boley proposed directed recursive labelnode hypergraph (DRLH) as a representation language for semantic network representations, capable of being specialised into natural language or other languages such as Lisp [11]. However, these methods did not develop graph models further to a direct hardware implementation, or present operational methods for database-level processing.
@@
Current GDB scalability challenges call for the development of specialised hardware accelerators that: a) can “natively” store information in graph format and b) can use in- and/or near-memory computation to perform graph-oriented operations on said data in a massively parallel fashion. In other words, we need to develop a data structure tailor-made for GDBs and an accompanying hardware that will store and operate on it. By processing graph data on such dedicated hardware, storage efficiency enhancement and direct real-time manipulations become feasible, leading to higher operation speed.
@@
In this work we propose Views, a graph database model tailored for use in hardware accelerators. First, we show how at its foundation it refactors descriptions of directed and labelled graphs as linked lists for better uniformity and symmetry in the data structure. Then we proceed to specify the exact data organisation structure that enables a piece of physical hardware to implement the GDB model and show both how it can represent a labelled graph and resolve semantic queries. Finally, we provide a few toy examples of how this GDB model could be used to perform simple “cognitive” tasks on a small dataset consisting of letters and strings.
@@
This paper is organised as follows: Section 2 describes the proposed Views GDB model in terms of its data encoding and structure, its equivalence to conventional graph representations, and its instantiation on graph database hardware. Section 3 presents the hardware implementation methodology, outlining 2 mapping schemes and Associative Chip Architecture (ASOCA) project for hardware acceleration. Section 4 demonstrates the operation on Views through examples in semantic reasoning and a Copycat-inspired cognitive processing application [12]. Section 5 provides a high-level discussion on the relationship of Views to other representations and potential extensions.
@@ 2 The Proposed Views Model
The proposed Views GDB model has been designed to represent graphs with the following features: (1) Directedness: vertex pairs are connected by directed edges (or “arcs”), (2) Labellability: edges and vertices can both be “labelled”, i.e., we can attach further properties to them and (3) Recursive properties: properties themselves can have properties and so on ad infinitum, rendering the capability of representing complex, nested relationships. We call these directed recursively labelled graphs, a term inspired by Boley’s DRLH [11] but restricting the underlying structure to graphs rather than hypergraphs, and any mention of “graph” in the rest of this paper will refer to a DRLG unless otherwise stated. Notably, both arbitrarily connected directed graphs and directed acyclic graphs are subsets of this definition.
We shall now introduce the basic data structure of the proposed GDB model and then explain its ability to support infinitely recursive labellability (i.e. infinitely recursive properties) as well as its mapping to hardware. Finally, the building and organisation of a Views-based GDB is presented.
@@
We begin construction of the Views data structure by considering a simple, non-trivial DRLG graph as illustrated in Figure 1a, which consists of two vertices (denoted in non-bevelled rectangles) and a single, labelled, directed edge connecting them. We call this a “half-$K_{2}$” digraph (to distinguish it from the full $K_{2}$ digraph which includes the reciprocal edge) and treat it as the “unit of relational information” within a graph: repeated instantiation of such ternary relationships can be used to create any DRLG graph. In the semantic web context, this corresponds to a Resource Description Framework (RDF) triple, and repeated instantiation constructs an RDF graph [13, 14]. Within the form of RDF triple, ¡subject-predicate-object¿, subject corresponds to the “source” vertex, predicate to the labelled edge, and object to the “destination” vertex. Finally, note that the terms “source”, “edge” and “destination” are used in a “de re” fashion and refer to the objects in question directly: we have not yet made use of labellability so far.
Views represents the conjugate ternary relationship using a “source-centred”, numerical structure that we shall call the Views triplet, meaning that edges and destination vertices are treated as equivalent entities (i.e., we make no distinction between them), both of which relate to the source vertex in the same fashion (as indicated by the “T” shape differing it from an RDF triple). This is illustrated in Figure 1b, where the source vertex “points to” both the edge and the destination vertex. Note that bevelled rectangles in the discourse differ from conventional vertices and they refer to physical memory entries as numbers or pointers, and that the arrows in this figure denote the concept of “owns”, i.e., the source vertex “owns” the marked destination vertex and edge. Thus, the pair of vertices and connecting edge have been mapped onto a set of 3 numbers that can be stored in a physical memory. So long as there is a way to track the association between the members of the Views triplet, we can represent a triplet in physical memory (more details in subsequent sections).
@@ 2.2 From Triplet to “Simple” Vertex-labelled GDB
With the linknode partly defined we can now add multiple linknodes with the same head ID in order to add further properties to some source vertex, each corresponds to an outgoing edge with its label and destination vertex of the source vertex in a graph. The collection of all linknodes “belonging to” some object X is called the chain of X. What all linknodes in a chain have in common is that they share head ID. The next pointer (at the bottom of Figure 2a) is added to allow the GDB to “traverse” (sequentially traverse) a chain, even if its linknodes are not stored sequentially in memory (i.e., the chain is “fragmented”). Starting from the head of a chain, next points to another linknode belonging to the same source vertex, and effectively turns a chain into a traversable linked list. Thus, an object with $N$ properties will feature $N$ linknodes arranged into a linked list. The objective here is to allow a near-memory processor to autonomously “discover” all linknodes attached to a particular source vertex/head ID without having to query the entire memory (e.g., via broadcast), which could be extremely large. Compare the cost of energising 32 billion memory entries to check if they are of interest, to following a couple of hundred linknodes by “hopping” from one to the other using the next pointer. Thus, the implementation of a next pointer is very much rooted in hardware considerations. We have now explained the entire structure of the proto-linknode.
@@ 2.3 The full Views Linknode: Adding Labellability
We now extend our proto-linknode structure to support recursive properties and become traversable in an elegant way. This upgrades the basic structure described in Figure 1a into what we see in Figure 4a: the full Views linknode data structure. Internally, a linknode is formatted as follows: [“source vertex”, “edge”, “edge properties”, “destination vertex”, “destination vertex properties”, “next linknode”], abbreviated to: [head ID, primID1, prop1, primID2, prop2, next], where primID stands for “primary ID” and “prop” for “properties”, illustrated in Figure 4a. The corresonding structure of headnode format is illustrated in Figure 4b.
In the previous section, we have covered the “labelling” process for source vertices, but edges and destination vertices have not yet been “labelled”. To do so, we furnish the edge and destination vertex entities with relationship-specific properties, illustrated in Figure 4a at the lower left and lower right (prop1 and prop2), to enable recursive labellability. As before, these are also numbers stored in a physical memory and each number corresponds to an address in memory, i.e., they are pointers. They point to linknodes that elaborate on the properties of primID1 and primID2 within the context of the linknode’s triplet, i.e., within the context of the relationship as a whole. The “context-dependent” versus “context-free” distinction is critical: In abstract terms it denotes the difference between “a property of object X” and “a property of object X when it relates to Y via Z”.
@@
Illustrative differences can be found between Figure 5b and Figure 5c, both of which describe the nested relationships in Figure 5a: it is up to the database schema whether to treat word meanings as context-dependent properties or as separate linknodes of their own. The linknode on the left (shrunk to a small orange circle for brevity) tells us that “cats belong to the family of Felidae”. In this particular example the “family” primID points to the head of the chain that contains information on what a “family” is, independent of context. Following the link and inspecting the chain we find that this particular “family” chain refers to the taxonomical context and is itself a noun. This does not depend in any way on the fact that cats belong to the family of Felidae, and in fact can be pointed to by multiple primdIDs scattered throughout the database.
@@ 3 Hardware Implementation
3.1 Mapping linknodes to physical arrays: the CNSM and Normalised mappings
@@ 3.2 The ASOCA implementation
Our Associative Chip Architecture (ASOCA) aims to turn the ideas behind Views into a series of hardware accelerators for GDB operations. We started by implementing a memory array; the Associative Memory Chip I (ASOCA1). Using beyond-CMOS [17] memory cells as the basic bit-storage unit, it is a dually-addressable memory (DAM) array capable of storing any of the pointers in the CNSM and Normalised schemes shown previously (i.e., any line in Table 1 and Table 2) and promises good power performance (further details in [18, 19]). ASOCA1 arrays are designed to hold $\times 64$-bit pointer entries each and for the purposes of this paper, they can be considered the “unit storage array”.
@@
Here is a basic example of how the ISA can be used to answer simple queries: after programming a complete database into the hardware with PROGs, massively parallel CARs on N1-type arrays can be used to quickly find the addresses of all linknodes belonging to a specific headnode (i.e., they can “highlight” a complete chain). This can be used directly for queries of the sort: “Fetch all information directly associated with Tom Hanks.” – note the use of “directly” to denote that we are not looking for information about Tom Hanks available in other chains, i.e. cases where thinking of a film or some other concept would “make us think of” Tom Hanks.

### v2

@@ Abstract
The graph database (GDB) is an increasingly common storage model for data involving relationships between entries. Beyond its widespread usage in database industries, the advantages of GDBs indicate a strong potential in constructing symbolic artificial intelligences and retrieval-augmented generation (RAG), where knowledge of data inter-relationships takes a critical role in implementation. However, current GDB models are not optimised for hardware acceleration, leading to bottlenecks in storage capacity and computational efficiency. In this paper, we propose a hardware-friendly GDB model, called Views. We show its data structure and organisation tailored for efficient storage and retrieval of graph data and demonstrate its functional equivalence and storage performance advantage compared to represent traditional graph representations. We further demonstrate its symbolic processing abilities in semantic reasoning and cognitive modelling with practical examples and provide a short perspective on future developments.
@@
Keywords – graph database, knowledge representation, near-memory computing, semantic processing
@@ 1 Introduction
Widespread usage of GDB is seen in knowledge representations, particularly in implementing knowledge graphs – most commonly instantiated as Resource Description Framework (RDF) triplestores or label property graphs[1]. RDF represents knowledge as “subject-predicate-object” triples with optional reification and named graphs for a clean logical footing [2, 3]. LPGs, on the other hand, attach labels and properties directly to vertices and edges for application-centric schemas and analytics. Across both families, the explicitness of knowledge graph semantics shows promise in symbolic cognitive computing by exposing symbols and relations to reasoning procedures [4, 5]. Knowledge graphs are also increasingly important as knowledge bases for large language models in RAG pipelines to improve their accuracy and reduce hallucinations [6]. These then call for more efficient graph data storage and processing methods to handle the ever-growing size and complexity of knowledge graphs.
@@
Operationally, querying is the critical workload in GDBs. For example, the query of “find me all films which Tom Hanks has acted in” followed by “who is Tom Hanks”, as anyone might do in any film database. Multiple query languages and formal frameworks have been proposed over the years for expressive graph pattern matching and path traversal in RDF and LPGs, such as GQL, SPARQL, Cypher, etc. However, these high-level languages are built upon but not necessarily optimised for existing computing facilities [7, 8]. To address these problems, contemporary solutions such as GPU clusters can alleviate some GDB workload throughput limits, although the high cost of power and deployment complexity constrain their widespread adoption [9].
As a result, contemporary GDB models and frameworks face significant challenges. In symbol-heavy applications like knowledge graphs, where content search operations dominate, this manifests as poor cache behaviour and bandwidth pressure. Heterogeneity of data types and cross-database operability further present a bottleneck in efficient processing and scaling [10, 11], for example in the domain of biological science[12]. Memory and schema optimisations are regarded as promising ways to enhance storage efficiency and query performance and to alleviate such issues [12, 13, 14], while practical hardware implementation considerations are still overlooked.
@@
Now we are going to talk about what to do in cases of data of unusual complexity. Historically, recursive labelling in GDB models has been proposed as a method to represent complex relationships in graphs, especially in semantic contexts. Pratt developed the initial hierarchical graph model of recursive labelling for semantics of programming languages [15], and Boley proposed directed recursive labelnode hypergraph (DRLH) as a representation language for semantic network representations, capable of being specialised into natural language or other languages such as Lisp [16]. However, these methods are not directly applicable to modern GDBs, due to the lack of support for graph operation practices and operational methods for database-level processing. They did not develop graph models further to a direct hardware implementation either.
@@
As an overview, current GDB scalability challenges call for the development of specialised hardware accelerators that: a) can “natively” store information in graph format and b) can use in- and/or near-memory computation to perform graph-oriented operations on said data in a massively parallel fashion. In other words, we need to develop a data structure tailor-made for GDBs and an accompanying hardware that will store and operate on it. By processing graph data on such dedicated hardware, storage efficiency enhancement and direct real-time manipulations become feasible, leading to higher operation speed.
@@
In this work, we propose Views, a graph database model tailored for use in hardware accelerators. First, we show how at its foundation it refactors descriptions of directed and labelled graphs as linked lists for better uniformity and symmetry in the data structure. Then we proceed to specify the exact data organisation structure that enables a piece of physical hardware to implement the GDB model and show both how it can represent a labelled graph and resolve semantic queries. Finally, we provide a few toy examples of how this GDB model could be used to perform simple “cognitive” tasks on a small dataset consisting of letters and strings.
@@
This paper is organised as follows: Section 2 describes the proposed Views GDB model in terms of its data encoding and structure, its equivalence to conventional graph representations, and its instantiation on graph database hardware. Section 3 presents the hardware implementation methodology with a comparison between existing graph database implementations, outlining 2 mapping schemes and Associative Chip Architecture (ASOCA) project for hardware acceleration. Section 4 demonstrates the operation on Views through examples in semantic reasoning and a Copycat-inspired cognitive processing application [17]. Section 5 provides a high-level discussion on the relationship of Views to other representations and potential extensions.
@@ 2 The Proposed Views Model
The proposed Views GDB model has been designed to represent graphs with the following features: (1) Directedness: vertex pairs are connected by directed edges (or “arcs”), (2) Labellability: edges and vertices can both be “labelled”, i.e., we can attach further properties to them and (3) Recursive properties: properties themselves can have properties and so on ad infinitum, rendering the capability of representing complex, nested relationships. We call these directed recursively labelled graphs, a term inspired by Boley’s DRLH [16] but restricting the underlying structure to graphs rather than hypergraphs, and any mention of “graph” in the rest of this paper will refer to a DRLG unless otherwise stated.
We shall now introduce the basic data structure of the proposed GDB model and then explain its ability to support infinitely recursive labellability as well as its mapping to hardware. Finally, the building and organisation of a Views-based GDB is presented.
@@
We begin construction of the Views data structure by considering a simple, non-trivial DRLG as illustrated in Figure 1a, which consists of two vertices (denoted in non-bevelled rectangles) and a single, labelled, directed edge connecting them. We call this a “half-$K_{2}$” digraph (to distinguish it from the full $K_{2}$ digraph which includes the reciprocal edge) and treat it as the “unit of relational information” within a graph: repeated instantiation of such ternary relationships can be used to create any DRLG. In the semantic web context, this corresponds to a RDF triple, and repeated instantiation constructs an RDF graph [2, 3]. Within the form of RDF triple, ¡subject-predicate-object¿, subject corresponds to the “source” vertex, predicate to the labelled edge, and object to the “destination” vertex. Finally, note that the terms “source”, “edge” and “destination” are used in a “de re” fashion and refer to the objects in question directly: we have not yet made use of labellability so far.
Views represents the conjugate ternary relationship using a “source-centred”, numerical structure that we shall call the Views triplet, meaning that edges and destination vertices are treated as equivalent entities (i.e., we make no distinction between them), both of which relate to the source vertex in the same fashion (as indicated by the “-t’ suffix’s shape). This is illustrated in Figure 1b, where the source vertex “points to” both the edge and the destination vertex. Note that bevelled rectangles in the discourse differ from conventional vertices and they refer to physical memory entries as numbers or pointers, and that the arrows in this figure denote the concept of “owns”, i.e., the source vertex “owns” the marked destination vertex and edge. Thus, the pair of vertices and connecting edge have been mapped onto a set of 3 numbers that can be stored in a physical memory. So long as there is a way to track the association between the members of the Views triplet, we can represent a triplet in physical memory (more details in subsequent sections).
@@ 2.2 From Triplet to “Simple” Vertex-labelled GDB
With the linknode partly defined we can now add multiple linknodes with the same head ID in order to add further properties to some source vertex; each corresponds to an outgoing edge with its label and destination vertex of the source vertex in a graph. The collection of all linknodes “belonging to” some object X is called the chain of X. What all linknodes in a chain have in common is that they share head ID. The next pointer (at the bottom of Figure 2a) is added to allow the GDB to “traverse” (sequentially traverse) a chain, even if its linknodes are not stored sequentially in memory (i.e., the chain is “fragmented”). Starting from the head of a chain, next points to another linknode belonging to the same source vertex, and effectively turns a chain into a traversable linked list. Thus, an object with $N$ properties will feature $N$ linknodes arranged into a linked list. The objective here is to allow a near-memory processor to autonomously “discover” all linknodes attached to a particular source vertex/head ID without having to query the entire memory (e.g., via broadcast), which could be extremely large. Compare the cost of energising 32 billion memory entries to check if they are of interest, to following a couple of hundred linknodes by “hopping” from one to the other using the next pointer. Thus, the implementation of a next pointer is very much rooted in hardware considerations. We have now explained the entire structure of the proto-linknode.
@@ 2.3 The full Views Linknode: Adding Labellability
We now extend our proto-linknode structure to support recursive properties and become traversable in an elegant way. This upgrades the basic structure described in Figure 1a into what we see in Figure 4a: the full Views linknode data structure. Internally, a linknode is formatted as follows: [“source vertex”, “edge”, “edge properties”, “destination vertex”, “destination vertex properties”, “next linknode”], abbreviated to: [head ID, primID1, prop1, primID2, prop2, next], where primID stands for “primary ID” and “prop” for “properties”, illustrated in Figure 4a. The corresponding structure of headnode format is illustrated in Figure 4b.
In the previous section, we have covered the “labelling” process for source vertices, but edges and destination vertices have not yet been “labelled”. To do so, we furnish the edge and destination vertex entities with relationship-specific properties, illustrated in Figure 4a at the lower left and lower right (prop1 and prop2), to enable recursive labellability. As before, these are also numbers stored in a physical memory and each number corresponds to an address in memory, i.e., they are pointers. They point to linknodes that elaborate on the properties of primID1 and primID2 within the context of the linknode’s triplet, i.e., within the context of the relationship as a whole. The “context-free” versus “context-dependent” distinction is critical: In abstract terms it denotes the difference between “a property of object X” and “a property of object X when it relates to Y via Z”.
@@
Illustrative differences can be found between Figure 5b and Figure 5c, both of which describe the nested relationships in Figure 5a: it is up to the database schema whether to treat word meanings as context-dependent properties or as separate linknodes of their own. The linknode on the left (shrunk to a small orange circle for brevity) tells us that “cats belong to the family of Felidae”. In this particular example the “family” primID points to the head of the chain that contains information on what a “family” is, independent of context. Following the link and inspecting the chain we find that this particular “family” chain refers to the taxonomical context and is itself a noun. This does not depend in any way on the fact that cats belong to the family of Felidae, and in fact can be pointed to by multiple primIDs scattered throughout the database.
@@ 3 Hardware Implementation
3.1 Mapping linknodes to physical arrays: Two Allocations
@@ 3.2 The ASOCA implementation
Our Associative Chip Architecture (ASOCA) aims to turn the ideas behind Views into a series of hardware accelerators for GDB operations. We started by implementing a memory array; the Associative Memory Chip I (ASOCA1). Using beyond-CMOS [21] memory cells as the basic bit-storage unit, it is a dually-addressable memory (DAM) array capable of storing any of the pointers in the CNSM and Normalised schemes shown previously (i.e., any line in Table 1 and Table 2) and promises good power performance (further details in [22, 23]). ASOCA1 arrays are designed to hold $64\times 64$-bit pointer entries each and for the purposes of this paper, they can be considered the “unit storage array”.
@@
Here is a basic example of how the ISA can be used to answer simple queries: after programming a complete database into the hardware with PROGs, massively parallel CARs on N1-type arrays can be used to quickly find the addresses of all linknodes belonging to a specific headnode (i.e., they can “highlight” a complete chain). This can be used directly for queries of the sort: “Fetch all information directly associated with Tom Hanks.” Note the use of “directly” to denote that we are not looking for information about Tom Hanks available in other chains, i.e. cases where thinking of a film or some other concept would “make us think of” Tom Hanks.
@@
The ISA under discussion is natively powered by Views to support the scalability and heterogeneity of data storage, while associative search ability from hardware memory arrays backs efficient data retrieval from Views-based GDBs. It denotes the strong synergy between the Views data structure and the ASOCA hardware architecture. Beyond advanced graph operations such as subgraph matching and graph traversal, this co-design methodology provides the feasibility for the dedicated hardware acceleration of optimised graph algorithms by utilising both the GDB model and the hardware architecture [25, 26].
3.3 Hardware Storage Comparisons
We compare the storage performance of a Views-based system against 4 representative GDB systems: Neo4j (LPG), Memgraph (LPG), Apache Jena TDB2(RDF) and Blazegraph (RDF). The benchmark reproduces the minimal dataset of Figure 7 in each implementation and in Views, and then measures total on-disk footprint.
Within LPG implementations, Neo4j uses native on-disk graph storage across multiple files, often fixed-size. Its disk usage is calculated from entity/property counts and documented per-record overheads [27]. Memgraph is primarily configured as an in-memory graph database, and we report its write-ahead log for disk usage.
Conventional RDF stores serialise datasets into plain formats for on-disk storage [14, 11], while both Blazegraph and Apache Jena TDB2 back indexes in B+ tree files. Blazegraph keeps a single journal file, and we extract its disk usage from internal statistics. TDB2 employs multiple files for indexing and doesn’t scale well with smaller datasets; we report the decompressed size of its N-Quads snapshot instead for comparison.
For Views, we implement the CNSM allocation on one ASOCA2 supercluster and measure the occupied memory after programming the dataset. Oversized strings are outsourced and indexed by primIDs; their sizes are included.
For reference purpose, we also record here each database’s dump size where possible (e.g., a snapshot or backup). For Views, this is the total disk usage with entry addresses included. The results are summarised in Table 3 and Figure 9.
Table 3: Storage Performance Comparison across Views, RDF and LPG Implementations
Database Disk usage (Bytes) Entity number Dump size (Bytes)
Neo4j 1554 24 7521
Memgraph 7362 24 446 (compressed)
Blazegraph 11697 39
Apache Jena TDB2 4014 39 540 (compressed)
Views 685 19 756
Figure 9: Storage efficiency comparison among different graph database implementations: (a) Disk usage in Bytes; (b) Number of entities stored.
Views stores relationships as directly addressable linknodes with no record headers or secondary indexes, which removes per-entity overhead as seen in RDF or LPGs. Second, pointer-based sharing avoids duplication: Views collapses a triple into one linknode plus reusable headnodes, while labels and objects are all referenced via headnodes. Last but not least, context-specific information is stored in sub-chains to enable direct hardware support for data locality and performance, rather than separate triples or property rows.
As a result, Views’ linknode data structure together with the matching hardware implementation means the model keeps bytes-per-relationship essentially flat, preserving its significant advantage in storage efficiency, retrieval performance and scalability. Still, note that some database implementations are affected by their on-disk storage format and page alignment. Optimisations, e.g., compression or index layout choices, can improve the numbers of existing GDB implementations quoted in Table 3. Views does not seek to entirely replace them, but instead offers strong compatibility with them as the underlying GDB model.